OpenStudy (anonymous):

A publisher for a promising new novel figures fixed costs (overhead, advances, promotion, copy editing, typesetting) at $55,000, and variable costs (printing, paper, binding, shipping) at $1.60 for each book produced. If the book is sold to distibutors for $11 each, how many must be produced and sold for the publisher to break even?

OpenStudy (anonymous):

Cost equations c(x) = 1.60x + 55000 Revenue equation r(x) = 11x : Break even is when Revenue = cost, therefore 11x = 1.60x + 55000 11x - 1.6x = 55000 9.4x = 55000 x = x ~ 5851 books ; : Prove this to your self Rev: 11(5851) = $64,361.00 Cost: 1.6(5851) + 55000 = $64,361.60 (slight loss we can sell a portion of a book)